In a class of 40 students, 1/5 of the total number of students like to eat rice only, 2/5 of the total number of students like to eat chapati only and the remaining students like to eat both. What fraction of the total number of students like to eat both?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Number of students in a class = 40 students
1/5 of the total number of students like to eat rice only = 1/5 × 40

=1 × 8

8 students
2/5 of the total number of students like to eat chapati only = 2/5 × 40
=2 × 8
= 16 students

Then,
Number of students like to eat both = 40–(8 + 16)
= 40–24
= 16 students
Fraction of the total number of students like to eat both = 16/40
= 2/5

Thanks